OpenStudy (anonymous):

2 √128 - √18 + 4 √32 I NEED HELP PLEASE!!!!! ITS AN EMERGENCY !!!! STEP BY STEP PLEASE!!!!!

OpenStudy (anonymous):

\[\sqrt{128}=\sqrt{16*4*2}=4*2\sqrt{2}=8\sqrt{2}\] \[\sqrt{18}=\sqrt{9*3}=3*\sqrt{3}\] \[\sqrt{32}=\sqrt{16*2}=4\sqrt{2}\] now, combine like terms...

OpenStudy (anonymous):

mistake...sorry.. \[\sqrt{18}=\sqrt{9*2}=3\sqrt{2}\] your answer: \[16\sqrt{2}-3\sqrt{2}+16\sqrt{2}=29\sqrt{2}\]